A man can jump over b=4m wide trench on earth. If mean density of an imaginary planet is twice thatof the earth, calculate its maximum possible radius so that he may escape from it by jumping. Givenradius of earth = 6400 km.

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
From the given data,
Mean density of planet = 2 density of earth
Mass / volume = 2 mass of earth / R^3
1/R^3 = 8 / (6400)^3
1/R = 2/6400
R = 3200m.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)